how much water should be added to 414 ounces of a 21% alcohol solution to get an 18% alcohol solution
i set it up this way---.21(414)+W=.18(414 + W) but i came up with a neg answer
86.94 + W = 74.52 + .18W
12.42 = -.82W or .82W = -12.42 === -.1546
I don't think the answer should be negative.

Answer by mananth(16946)

how much water should be added to 414 ounces of a 21% alcohol solution to get an 18% alcohol solution
how much water ------------0% alcohol-----------x ounces
alcohol ------------------ 21% alcohol----------414 ounces
alcohol------------------- 18% alcohol----------(x+414)ounces
..
concentration equation
...
0+0.21*414=0.18(x+414)
0+86.94=0.18x+74.52
0.18x=86.94-74.52
0.18x=12.42
x=12.42/0.18
x= 69 ounces of water
..
m.ananth@hotmail.ca